### v2

# Hopf orbits and the first ECH capacity

Abstract.
We consider dynamically convex (star-shaped) domains in a symplectic vector space of dimension $4$. For such a domain, a “Hopf orbit” is a closed characteristic in the boundary which is unknotted and has self-linking number $-1$. We show that the minimum action among Hopf orbits exists (without any nondegeneracy hypothesis) and defines a symplectic capacity for dynamically convex domains in four dimensions. We further show that this capacity agrees with the first ECH capacity for dynamically convex domains. Combined with a result of Edtmair, this implies that for dynamically convex domains in four dimensions, the first ECH capacity agrees with the cylinder capacity.
1. Introduction
1.1. Symplectic capacities and the Viterbo conjecture
Recall that if $(X_{1},\omega_{1})$ and $(X_{2},\omega_{2})$ are symplectic manifolds of the same dimension, a symplectic embedding of $X_{1}$ into $X_{2}$ is a smooth embedding $\varphi:X_{1}\hookrightarrow X_{2}$ such that $\varphi^{*}\omega_{2}=\omega_{1}$. In this paper, the symplectic manifolds that we consider will usually be domains [^1] Here we define a “domain” in $\mathbb{R}^{2n}$ to be the closure of a nonempty open set. in $\mathbb{R}^{2n}=\mathbb{C}^{n}$ with smooth boundary, usually with $n=2$, which we always equip with the restriction of the standard symplectic form
$\omega_{0}=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}.$
In particular, define the ball
$B^{2n}(r)=\{z\in\mathbb{C}^{n}\mid\pi|z|^{2}\leq r\}$
and the cylinder
$Z^{2n}(R)=\{z\in\mathbb{C}^{n}\mid\pi|z_{1}|^{2}\leq R\}.$
In 1985 Gromov proved the celebrated nonsqueezing theorem [gromov], asserting that there exists a symplectic embedding $B^{2n}(r)\hookrightarrow Z^{2n}(R)$ if and only if $r\leq R$. Since then, symplectic embeddings have been extensively studied, see e.g. [35].
One of the main tools in this study are symplectic capacities, see e.g. the survey [7]. Definitions of this term in the literature vary, and we will use the following:
Definition 1.1.
A symplectic capacity defined in dimension $2n$ is a function $c$ from some set of $2n$-dimensional symplectic manifolds (including at least convex domains in $\mathbb{R}^{2n}$) to $[0,\infty]$, such that:
  • (Monotonicity) If $(X_{1},\omega_{1})$ and $(X_{2},\omega_{2})$ are symplectic manifolds of dimension $2n$ for which $c$ is defined, and if there exists a symplectic embedding $X_{1}\hookrightarrow X_{2}$, then
    (1.1) $c(X_{1})\leq c(X_{2}).$
  • (Conformality) If $r>0$ and if $c$ is defined on $(X,\omega)$, then
    (1.2) $c(X,r\omega)=rc(X,\omega).$
    In particular, if $X$ is a domain in $\mathbb{R}^{2n}$, this means that $c(rX)=r^{2}c(X)$.
We say that a symplectic capacity $c$ is normalized if in addition
$c(B^{2n}(r))=c(Z^{2n}(r))=r.$
The simplest examples of normalized symplectic capacities are the ball capacity $c_{B}$ and the cylinder capacity $c_{Z}$. If $\operatorname{dim}(X)=2n$, then these are defined by
$c_{B}(X,\omega)$ $=\sup\{r\mid\mbox{$\exists$ symplectic embedding $B^{2n}(r)\hookrightarrow(X,\omega)$}\},$
(1.3) $c_{Z}(X,\omega)$ $=\inf\{R\mid\mbox{$\exists$ symplectic embedding $(X,\omega)\hookrightarrow Z^{2n}(R)$}\}.$
The statements that $c_{B}$ and $c_{Z}$ are normalized symplectic capacities are equivalent to Gromov’s nonsqueezing theorem. In addition, it follows from the definitions that if $c$ is a normalized symplectic capacity defined on $(X,\omega)$, then
(1.4) $c_{B}(X,\omega)\leq c(X,\omega)\leq c_{Z}(X,\omega).$
The following conjecture, possibly due to Hofer, has apparently been folklore since the 1990s:
Conjecture 1.2 (“strong Viterbo conjecture”).
All normalized symplectic capacities agree on compact convex domains in $\mathbb{R}^{2n}$.
By (1.4), this conjecture is equivalent to the statement that if $X$ is a compact convex domain in $\mathbb{R}^{2n}$, then $c_{B}(X)=c_{Z}(X)$. The conjecture is known for certain special classes of convex domains, for example for $S^{1}$-invariant convex domains by Ostrover, as reviewed in [GHR, Prop. 1.4], for convex domains in a $C^{3}$ neighborhood of a round ball in $\mathbb{R}^{4}$ by Edtmair [gss_edtmair], and more recently for domains in a $C^{2}$ neighborhood of a round ball in $\mathbb{R}^{2n}$ in [1]. In addition, some normalized symplectic capacities are known to agree on convex domains, as we discuss below.
Various symplectic capacities are defined in terms of Reeb dynamics and contact homology. To say more about this, we need the following definitions.
Definition 1.3.
A star-shaped domain is a compact domain $X\subset\mathbb{R}^{2n}$ with smooth connected boundary such that $0\in\operatorname{int}(X)$ and $\partial X$ is transverse to the radial vector field.
If $X$ is a star-shaped domain as above, then the standard Liouville form on $\mathbb{R}^{2n}$ defined by
$\lambda_{0}=\frac{1}{2}\sum_{i=1}^{n}\left(x_{i}\,dy_{i}-y_{i}\,dx_{i}\right)$
restricts to a contact form $\lambda$ on $\partial X$. The contact form $\lambda$ determines the Reeb vector field $R_{\lambda}$ characterized by $d\lambda(R_{\lambda},\cdot)=0$ and $\lambda(R_{\lambda})=1$. A Reeb orbit is a periodic orbit of $R_{\lambda}$, i.e. a map $\gamma:\mathbb{R}/T\mathbb{Z}\to\partial X$ for some $T>0$, modulo precomposition with translations of $\mathbb{R}/T\mathbb{Z}$, such that $\gamma^{\prime}(t)=R_{\lambda}(\gamma(t))$ for all $t$. The Reeb orbit $\gamma$ is simple if the map $\gamma$ is an embedding. The symplectic action $\mathcal{A}(\gamma)>0$ is defined to be the period $T$. By a theorem of Rabinowitz [34], at least one Reeb orbit exists. Let $\mathcal{P}(\partial X)$ denote the set of Reeb orbits in $\partial X$, and define
$\mathcal{A}_{\operatorname{\min}}(X)=\min_{\gamma\in\mathcal{P}(\partial X)}\mathcal{A}(\gamma).$
The Ekeland-Hofer-Zehnder capacity is a normalized symplectic capacity $c_{\operatorname{EHZ}}$ which is defined for convex domains in $\mathbb{R}^{2n}$ as follows. If $X$ is also a star-shaped domain as above, then
(1.5) $c_{\operatorname{EHZ}}(X)=\mathcal{A}_{\operatorname{min}}(X).$
It is a nontrivial fact that $\mathcal{A}_{\operatorname{min}}$ is monotone under symplectic embeddings, and then it extends by $C^{0}$ continuity to all convex domains [^2] The history is that the previously defined Ekeland-Hofer and Hofer-Zehnder capacities were shown to agree with the minimal action $\mathcal{A}_{\operatorname{min}}$ for convex star-shaped domains; see [GHR, Thm. 1.12]..
The ECH capacities are a sequence of symplectic capacities $\{c_{k}^{\operatorname{ECH}}\}_{\geq 0}$ defined in four dimensions [hut_qech]. If $(X,\omega)$ is a four-dimensional symplectic manifold, then
$0=c_{0}^{\operatorname{ECH}}(X,\omega)<c_{1}^{\operatorname{ECH}}(X,\omega)\leq c_{2}^{\operatorname{ECH}}(X,\omega)\leq\cdots\leq\infty.$
If $X\subset\mathbb{R}^{2n}$ is a star-shaped domain, then for each $k$ there is a finite set of Reeb orbits $\{\gamma_{i}\}$ on $\partial X$, selected homologically by embedded contact homology, such that
$c_{k}^{\operatorname{ECH}}(X)=\sum_{i}\mathcal{A}(\gamma_{i}).$
The first ECH capacity $c_{1}^{\operatorname{ECH}}$ is normalized, and we will review its definition for star-shaped domains in §3.2.
There are also “alternative ECH capacities” $c_{k}^{\operatorname{Alt}}$ defined by a more elementary max-min formula in [28], which satisfy the same basic properties as those of the ECH capacities listed above and agree with them on basic examples. If $X$ is a star-shaped domain, then by [28, Thm. 12] we have
$c_{k}^{\operatorname{Alt}}(X)\leq c_{k}^{\operatorname{ECH}}(X).$
Some additional sequences of symplectic capacities, defined for star-shaped domains in $\mathbb{R}^{2n}$, are the Ekeland-Hofer capacities $\{c_{k}^{\operatorname{EH}}\}_{k\geq 1}$ defined in [11] using variational methods, and the conjecturally equal capacities $\{c_{k}^{\operatorname{CH}}\}_{k\geq 1}$ defined in [13] using equivariant symplectic homology. The capacities $c_{1}^{\operatorname{EH}}$ and $c_{1}^{\operatorname{CH}}$ are normalized. As reviewed in [GHR, Thm. 1.12], if $X$ is a convex domain in $\mathbb{R}^{2n}$ then it is known that
$c_{\operatorname{EHZ}}(X)=c_{1}^{\operatorname{EH}}(X)=c_{1}^{\operatorname{CH}}(X).$
For a convex domain in $\mathbb{R}^{4}$ one can further show that the above capacities agree with $c_{1}^{\operatorname{Alt}}(X)$. However it is not known whether the latter agrees with $c_{1}^{\operatorname{ECH}}(X)$; see [28, Rem. 13].
Remark 1.4.
The original Viterbo conjecture from [viterbo] asserts that if $c$ is a normalized symplectic capacity, and if $X$ is a compact convex domain in $\mathbb{R}^{2n}$, then
(1.6) $c(X)^{n}\leq n!\operatorname{vol}(X),$
with equality if and only if $\operatorname{int}(X)$ is symplectomorphic to an open round ball. Note that the inequality (1.6) follows from the strong Viterbo conjecture. This is because if there exists a symplectic embedding $B^{2n}(r)\hookrightarrow X$, then
$\frac{r^{n}}{n!}=\operatorname{vol}(B^{2n}(r))\leq\operatorname{vol}(X),$
and it follows that $c_{B}(X)^{n}\leq n!\operatorname{vol}(X)$.
Remark 1.5.
The “weak Viterbo conjecture” asserts that if $X$ is a convex star-shaped domain, then $\mathcal{A}_{\operatorname{min}}(X)^{n}\leq n!\operatorname{vol}(X)$. This follows from the inequality in the Viterbo conjecture (1.6) and equation (1.5) since $c_{\operatorname{EHZ}}$ is a normalized symplectic capacity. The weak Viterbo conjecture is known to imply the Mahler conjecture in convex geometry [3]. By continuity, the weak Viterbo conjecture is equivalent to the statement that if $X$ is a convex polytope, then $c_{\operatorname{EHZ}}(X)^{n}\leq n!\operatorname{vol}(X)$. The latter conjecture has been tested numerically in [6, 15].
1.2. The Hopf capacity
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain and let $\lambda$ denote the contact form on $\partial X$. The contact structure $\xi=\operatorname{Ker}(\lambda)$ has, up to homotopy, a unique symplectic trivialization $\tau$ over $\partial X$. If $\gamma$ is a Reeb orbit, then the linearized Reeb flow along $\gamma$ has a well-defined rotation number with respect to $\tau$, which we denote by $\rho(\gamma)\in\mathbb{R}$; see [27, §1.2] for details.
Definition 1.6.
A star-shaped domain $X\subset\mathbb{R}^{4}$ is dynamically convex if every Reeb orbit $\gamma$ in $\partial X$ has rotation number $\rho(\gamma)>1$.
Remark 1.7.
The definition of dynamical convexity is more commonly stated as the condition that every Reeb orbit $\gamma$ in $\partial X$ has Conley-Zehnder index $\operatorname{CZ}(\gamma)\geq 3$. Here if $\gamma$ is a nondegenerate Reeb orbit (see §2 for the definition of “nondegenerate”), then
(1.7) $\operatorname{CZ}(\gamma)=\lfloor\rho(\gamma)\rfloor+\lceil\rho(\gamma)\rceil.$
If $\gamma$ is degenerate, which in particular implies that $\rho(\gamma)$ is an integer, then there are different possible conventions for the Conley-Zehnder index. We will use the maximal lower semicontinuous extension of the nondegenerate Conley-Zehnder index, which in the degenerate case sometimes agrees with (1.7) and sometimes is $1$ less; see §2.2 for the precise definition. With this convention, $\rho(\gamma)>1$ is equivalent to $\operatorname{CZ}(\gamma)\geq 3$.
The notion of dynamical convexity was introduced by Hofer-Wysocki-Zehnder [convex], as a symplectomorphism-invariant alternative to the notion of convexity. It is shown in [convex] that every strictly convex star-shaped domain in $\mathbb{R}^{4}$ is dynamically convex. For a long time it was not known whether, conversely, every dynamically convex domain is symplectomorphic to a convex domain. It was recently shown by Chaidez-Edtmair [5] that the answer is no; see [9, 8] for further developments. In addition, it is shown in [ABHS_comp] that the conclusion of the weak Viterbo conjecture fails for some dynamically convex domains in $\mathbb{R}^{4}$. Nonetheless, the results that we will prove in this paper are valid for dynamically convex domains.
Recall that the self-linking number of a simple Reeb orbit in $\partial X$, or more generally any knot transverse to the contact structure $\xi$, is its linking number with a pushoff via the framing induced by the global trivialization $\tau$ of $\xi$.
Definition 1.8.
A Hopf orbit is a simple Reeb orbit $\gamma$ in $\partial X$ which is unknotted and has self-linking number $-1$. Let $\mathcal{P}_{\operatorname{Hopf}}(\partial X)$ denote the set of Hopf orbits in $\partial X$.
Remark 1.9.
It is shown in [convex] that if $X$ is dynamically convex, then the Reeb flow on $\partial X$ admits a disk-like global surface of section. This is an embedded disk $D\subset\partial X$ such that $\partial D$ is a simple Reeb orbit $\gamma$; the Reeb flow is transverse to $\mathrm{int}(D)$; and for every $x\in\partial X\setminus\gamma$, the Reeb trajectory through $x$ intersects $\mathrm{int}(D)$ forwards and backwards in time. It follows from these conditions that $\gamma$ is a Hopf orbit. Conversely, it is shown in [hryn_jsg] that if $X$ is dynamically convex, then every Hopf orbit in $\partial X$ bounds a disk-like global surface of section for the Reeb flow.
Definition 1.10.
If $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, define
$\mathcal{A}_{\rm Hopf}(X)=\inf_{\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)}\mathcal{A}(\gamma).$
We can use $\mathcal{A}_{\rm Hopf}$ in a standard way to define an invariant of symplectic 4-manifolds as follows.
Definition 1.11.
If $(X,\omega)$ is a symplectic four-manifold, define
(1.8) $c_{\rm Hopf}(X,\omega)\ =\ \sup\{\mathcal{A}_{\rm Hopf}(\Omega)\mid\Omega\text{ is dynamically convex and }\Omega\hookrightarrow(X,\omega)\}$
where the supremum is taken over all dynamically convex domains $\Omega\subset\mathbb{R}^{4}$ such that $\Omega$ symplectically embeds into $(X,\omega)$.
Our first result is the following.
Theorem 1.12.
The invariant $c_{\rm Hopf}$ is a normalized symplectic capacity. Moreover, if $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then
$c_{\rm Hopf}(X)=\mathcal{A}_{\rm Hopf}(X)=\mathcal{A}(\gamma)$
for some $\gamma\in\mathcal{P}_{\rm Hopf}(X)$ satisfying ${\rm CZ}(\gamma)=3$.
Let $X\subset\mathbb{R}^{4}$ be a dynamically convex domain. It follows from Theorem 1.12 and (1.4) that
$\mathcal{A}_{\rm Hopf}(X)\leq c_{Z}(X).$
On the other hand, Edtmair [gss_edtmair] showed that if there exists a Hopf orbit $\gamma$ with $\mathcal{A}(\gamma)=R$, then there exists a symplectic embedding $X\hookrightarrow Z(R)$. In particular, $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$. Combining this with the second part of Theorem 1.12, we get:
Corollary 1.13.
If $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then
(1.9) $c_{Z}(X)=\mathcal{A}_{\rm Hopf}(X),$
and the infimum in (1.3) is realized.
1.3. The first ECH capacity
Our second result is the following dynamical interpretation of the first ECH capacity.
Theorem 1.14.
If $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then
$c_{1}^{\rm ECH}(X)=\mathcal{A}_{\operatorname{Hopf}}(X).$
Remark 1.15.
One of the consequences of Theorems 1.12 and 1.14 and their proofs is that we can prove that $c_{1}^{\rm ECH}$ is a symplectic capacity for dynamically convex domains, without using Seiberg-Witten theory as in the proof of monotonicity of ECH capacities in [hut_qech].
Remark 1.16.
The first ECH capacity of a dynamically convex domain $X\subset\mathbb{R}^{4}$ is the first number in the ECH spectrum of $\partial X$ with the induced contact form. Shibata [36, 37] has independently established dynamical interpretations of the first number in the ECH spectrum, similar to Theorem 1.14 and with some similar arguments, for certain lens spaces with contact forms that are either strictly convex or dynamically convex and nondegenerate. Note that Theorem 1.14 does not make any nondegeneracy assumption.
To summarize the situation with the strong Viterbo conjecture, for a convex domain $X\subset\mathbb{R}^{4}$, we now know that the normalized symplectic capacities discussed above can be sorted in nondecreasing order as follows:
  • The ball capacity $c_{B}$.
  • The capacities $c_{\operatorname{EHZ}}=c_{1}^{\operatorname{EH}}=c_{1}^{\operatorname{CH}}=c_{1}^{\operatorname{Alt}}$.
  • The capacities $c_{\operatorname{Hopf}}=c_{1}^{\operatorname{ECH}}=c_{Z}$. (These equalities hold more generally in the dynamically convex case.)
Acknowledgments. U.H. acknowledges the support by DFG SFB/TRR 191 ‘Symplectic Structures in Geometry, Algebra and Dynamics’, Projektnummer 281071066-TRR 191. M.H. was partially supported by NSF grant DMS-2005437. V.G.B.R. is grateful for the hospitality of the Institute for Advanced Study, where part of this work was completed. V.G.B.R. was partially supported by NSF grant DMS-1926686, FAPERJ grant JCNE E-26/201.399/2021 and a Serrapilheira Institute grant.
2. The Hopf capacity
The goal of this section is to prove Theorem 1.12. Before stating the main lemma, we recall a few definitions from contact dynamics.
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain, and let $\lambda=\lambda_{0}|_{\partial X}$ be the induced contact for on $\partial X$. Then $d\lambda$ restricts to a symplectic structure on the contact structure $\xi=\operatorname{Ker}(\lambda)$. We denote by $\phi^{t}:\partial X\to\partial X$ the flow of the Reeb vector field $R_{\lambda}$. The derivative of the flow restricts to a symplectic linear map
$d\phi^{t}:(\xi_{p},d\lambda)\longrightarrow(\xi_{\phi^{t}(p)},d\lambda).$
Now let $\gamma:\mathbb{R}/T\mathbb{Z}\to\partial X$ be a Reeb orbit. The Reeb orbit $\gamma$ is nondegenerate if $1$ is not in the spectrum of
(2.1) $d\phi^{T}:{\xi}_{\gamma(0)}\longrightarrow{\xi}_{\gamma(0)}.$
We say that the domain $X$ (or the contact form on $\partial X$) is nondegenerate (resp. nondegenerate up to action $C$) if all Reeb orbits $\gamma\in\mathcal{P}(\partial X)$ (resp. all Reeb orbits with action less than $C$) are nondegenerate. Finally, we say that the domain $X$ (or the contact form on $\partial X$) is dynamically convex up to action $C$ if ${\rm CZ}(\gamma)\geq 3$ for all $\gamma\in\mathcal{P}(\partial X)$ with $\mathcal{A}(\gamma)<C$.
2.1. The main lemma and the proof of Theorem 1.12
The crucial step in the proof of Theorem 1.12 is the following result.
Lemma 2.1.
Let $\Omega_{1}$ and $\Omega_{2}$ be star-shaped domains in $\mathbb{R}^{4}$. Let $C>\mathcal{A}_{\rm Hopf}(\Omega_{2})$ and assume that $\Omega_{1},\Omega_{2}$ are, up to action $C$, nondegenerate and dynamically convex. Suppose that there exists a symplectic embedding $\varphi:\Omega_{1}\hookrightarrow\mathrm{int}(\Omega_{2})$. Then there exists a Hopf orbit $\gamma\in\mathcal{P}_{\rm Hopf}(\Omega_{1})$ such that ${\rm CZ}(\gamma)=3$ and $\mathcal{A}(\gamma)<\mathcal{A}_{\rm Hopf}(\Omega_{2})$. In particular, $\mathcal{A}_{\rm Hopf}(\Omega_{1})<\mathcal{A}_{\rm Hopf}(\Omega_{2})$.
Proof of Theorem 1.12, assuming Lemma 2.1..
We proceed in five steps.
Step 1
We first show that $c_{\operatorname{Hopf}}$ is a symplectic capacity.
It follows immediately from equation (1.8) that $c_{\operatorname{Hopf}}$ satisfies the Monotonicity property (1.1).
We next show that $c_{\operatorname{Hopf}}$ satisfies the Conformality property (1.2). By equation (1.8), it is enough to show that if $\Omega\subset\mathbb{R}^{4}$ is a dynamically convex domain and $r>0$, then
(2.2) $\mathcal{A}_{\operatorname{Hopf}}(r\Omega)=r^{2}\mathcal{A}_{\operatorname{Hopf}}(\Omega).$
Equation (2.2) holds because the scaling map $\Omega\to r\Omega$ preserves the direction of the Reeb vector field but multiplies the period of all Reeb orbits by $r^{2}$.
Step 2
We now set up some notation.
If $g:S^{3}\to\mathbb{R}^{>0}$ is a smooth function, define a star-shaped domain
$\Omega_{g}=\{z\in\mathbb{R}^{4}\mid|z|\leq g(z/|z|)\}.$
Every star-shaped domain in $\mathbb{R}^{4}$ can be uniquely described this way. There is a diffeomorphism $S^{3}\longrightarrow\partial\Omega_{g}$ sending $z\mapsto g(z)z$. This diffeomorphism pulls back the contact form on ${\lambda_{0}}|_{\partial X_{g}}$ to $g^{2}({\lambda_{0}}|_{S^{3}})$.
Step 3
We now show that if $X\subset\mathbb{R}^{4}$ is a dynamically convex domain which is nondegenerate up to action $C>\mathcal{A}_{\rm Hopf}(X)$, then
(2.3) $c_{\rm Hopf}(X)=\mathcal{A}_{\rm Hopf}(X).$
To prove (2.3), first note that it follows from (1.8) that $c_{\rm Hopf}(X)\geq\mathcal{A}_{\rm Hopf}(X)$. To prove the reverse inequality, suppose that $\Omega$ is dynamically convex and that there exists a symplectic embedding $\Omega\hookrightarrow X$. We need to show that
(2.4) $\mathcal{A}_{\operatorname{Hopf}}(\Omega)\leq\mathcal{A}_{\operatorname{Hopf}}(X).$
To do so, let $\varepsilon>0$ be given. Let $\lambda$ denote the contact form on $\partial\Omega$. By a standard genericity argument, there is a function
$f:\partial\Omega\longrightarrow(1,1+\varepsilon)$
such that the contact form $f^{2}\lambda$ is nondenegerate. By a compactness argument (compare Step 5 below), we can choose $f$ so that $f^{2}\lambda$ is also dynamically convex up to action $C$. By Step 2 we can write $\Omega=\Omega_{g}$ for some $g:S^{3}\to\mathbb{R}^{>0}$. We then have symplectic embeddings
$\Omega\hookrightarrow\Omega_{fg}\hookrightarrow(1+\varepsilon)X.$
Applying Lemma 2.1 twice and equation (2.2), we obtain
$\mathcal{A}_{\rm Hopf}(\Omega)<\mathcal{A}_{\rm Hopf}(\Omega_{fg})<\mathcal{A}_{\rm Hopf}((1+\varepsilon)X)=(1+\varepsilon)^{2}\mathcal{A}_{\rm Hopf}(X).$
Since $\varepsilon>0$ was arbitrary, the inequality (2.4) follows.
Step 4
We now show that the symplectic capacity $c_{\operatorname{Hopf}}$ is normalized. Given real numbers $a,b>0$, define the ellipsoid
$E(a,b)=\left\{z\in\mathbb{C}^{2}\;\bigg{|}\;\frac{\pi|z_{1}|^{2}}{a}+\frac{\pi|z_{2}|^{2}}{b}\leq 1\right\}.$
There are two simple Reeb orbits in $\partial E(a,b)$, given by the circles $(z_{2}=0)$ and $(z_{1}=0)$. These Reeb orbits have actions $a$ and $b$ respectively, and they are both Hopf orbits. If $a/b$ is rational then there are additional simple Reeb orbits, but these do not have minimal action. Using Step 3, we conclude that
(2.5) $c_{\operatorname{Hopf}}(E(a,b))=\mathcal{A}_{\operatorname{Hopf}}(E(a,b))=\min(a,b).$
In particular, since $E(r,r)=B^{4}(r)$, we obtain $c_{\operatorname{Hopf}}(B^{4}(r))=r$.
To prove that $c_{\operatorname{Hopf}}(Z(R))=R$, let $\Omega\subset\mathbb{R}^{4}$ be a dynamically convex domain and suppose that there exists a symplectic embedding $\varphi:\Omega\hookrightarrow Z(R)$. We need to show that $\mathcal{A}_{\operatorname{Hopf}}(\Omega)\leq R$. Given $\varepsilon>0$, there exists $C>>0$ such that $\varphi(\Omega)\subset E((1+\varepsilon)R,C)$. By Step 3, Monotonicity, and equation (2.5), it follows that
$\mathcal{A}_{\operatorname{Hopf}}(\Omega)=c_{\operatorname{Hopf}}(\Omega)\leq c_{\operatorname{Hopf}}(E((1+\varepsilon)R,C))\leq(1+\varepsilon)R.$
Since $\varepsilon>0$ was arbitrary, we conclude that $\mathcal{A}_{\operatorname{Hopf}}(\Omega)\leq R$ as desired.
Step 5
To complete the proof of Theorem 1.12, we need to show that if $X\subset\mathbb{R}^{4}$ is a dynamically convex domain, then there exists $\gamma\in\mathcal{P}_{\rm Hopf}(\partial X)$ with ${\rm CZ}(\gamma)=3$ such that $\mathcal{A}(\gamma)=\mathcal{A}_{\rm Hopf}(X)$.
Let $\{\gamma_{n}\}_{n=1,\ldots}$ be a sequence of Hopf orbits in $\mathcal{P}_{\rm Hopf}(\partial X)$ with $\mathcal{A}(\gamma_{n})<e\mathcal{A}_{\operatorname{Hopf}}(X)$ and $\lim_{n\to\infty}\mathcal{A}(\gamma_{n})=\mathcal{A}_{\operatorname{Hopf}}(X)$. We remark that by a compactness argument and the Arzeli-Ascoli theorem, one can pass to a subsequence so that $\gamma_{n}$ converges to a Reeb orbit with action $\mathcal{A}_{\operatorname{Hopf}}(X)$. However this not enough to prove the theorem, because the limiting Reeb orbit might not be simple or have Conley-Zehnder index 3. We will instead need to take a limit of a sequence of Reeb orbits of different contact forms.
Let $\lambda$ denote the contact form on $\partial X$. By Step 2, replacing $X$ with a different star-shaped domain is equivalent to multiplying $\lambda$ by a positive function. We can construct a sequence of functions $f_{n}:\partial X\to(0,1)$ with $f_{n}\to 0$ in $C^{\infty}$ such that:
  • The contact form $e^{f_{n}}\lambda$ is nondegenerate and dynamically convex up to action $e^{3}\mathcal{A}_{\operatorname{Hopf}}(X)$.
  • The Reeb vector field of $e^{f_{n}}\lambda$ is tangent to $\gamma_{n}$, so that the loop $\gamma_{n}$, suitably reparametrized, defines a Reeb orbit $\widetilde{\gamma}_{n}\in\mathcal{P}_{\operatorname{Hopf}}(e^{f_{n}}\lambda)$.
Note that
(2.6) $\mathcal{A}(\gamma_{n})<\mathcal{A}(\widetilde{\gamma}_{n})<e^{\max(f_{n})}\mathcal{A}(\gamma_{n})<e\mathcal{A}(\gamma_{n})<e^{2}\mathcal{A}_{\operatorname{Hopf}}(X).$
In particular,
(2.7) $\lim_{n\to\infty}\mathcal{A}(\widetilde{\gamma}_{n})=\mathcal{A}_{\operatorname{Hopf}}(X).$
Now for each $n$, choose a constant $\epsilon_{n}\in(0,\min(f_{n}))$. The contact form $e^{f_{n}-\epsilon_{n}}\lambda$ is nondegenerate and dynamically convex up to action $e^{2}\mathcal{A}_{\operatorname{Hopf}}(X)$, since it is a constant scaling of $e^{f_{n}}\lambda$. By Lemma 2.1 and (2.6), there is a Hopf orbit $\gamma_{n}^{\prime}\in\mathcal{P}_{\operatorname{Hopf}}(e^{f_{n}-\epsilon_{n}}\lambda)$ such that
(2.8) $\mathcal{A}(\gamma_{n}^{\prime})<\mathcal{A}(\widetilde{\gamma}_{n})$
and $\operatorname{CZ}(\gamma_{n}^{\prime})=3$.
Now by a compactness argument using the Arzela-Ascoli theorem, we can pass to a subsequence so that $\gamma_{n}^{\prime}$ converges to a Reeb orbit $\gamma\in\mathcal{P}(\partial X)$. By (2.7) and (2.8), we have $\mathcal{A}(\gamma)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$. So to complete the proof, we just need to show that $\gamma$ is a Hopf orbit and $\operatorname{CZ}(\gamma)=3$.
By dynamical convexity and the lower semicontinuity of the Conley-Zehnder index, we have ${\rm CZ}(\gamma)=3$. If $\gamma$ is a simple Reeb orbit, then $\gamma$ is a Hopf orbit and we are done. If $\gamma$ is not simple, then $\gamma$ is a multiple cover of some Reeb orbit $\gamma^{\prime}$, which by dynamical convexity has rotation number $\rho(\gamma^{\prime})>1$. It follows that the rotation number $\rho(\gamma)>2$, so ${\rm CZ}(\gamma)\geq 5$, which is a contradiction. ∎
We now prepare for the proof of Lemma 2.1, which occupies the rest of §2.
2.2. Reeb orbits, asymptotic operators, and Conley-Zehnder indices
Let $\lambda$ be a contact form on a 3-manifold $Y$. Let $J$ be a complex structure on the vector bundle $\xi=\ker(\lambda)$ which is compatible with $d\lambda$. Given a Reeb orbit $\gamma$ with period $T$, the asymptotic operator is the unbounded operator $L_{\gamma,J}$ on $L^{2}(\gamma(T\cdot)^{*}\xi)$ sending
$\eta\longmapsto-J\nabla_{t}\eta$
where $\nabla$ is the connection on $\gamma(T\cdot)^{*}\xi$ induced by the linearized Reeb flow. One can check that this operator is self-adjoint when $L^{2}(\gamma(T\cdot)^{*}\xi)$ is equipped with the inner-product induced by $(d\lambda,J)$.
[…]
$\operatorname{wind}_{\tau}(\nu)\in\mathbb{Z}.$
For every $k\in\mathbb{Z}$ there are precisely two eigenvalues counted with multiplicity satisfying $\operatorname{wind}_{\tau}=k$, and $\nu_{1}\leq\nu_{2}\Rightarrow\operatorname{wind}_{\tau}(\nu_{1})\leq\operatorname{wind}_{\tau}(\nu_{2})$. For $\delta\in\mathbb{R}$ we set
$\alpha^{<\delta}_{\tau}(\gamma)$ $=\max\ \{\operatorname{wind}_{\tau}(\nu)\mid\nu\ \text{eigenvalue},\ \nu<\delta\}$
$\alpha^{\geq\delta}_{\tau}(\gamma)$ $=\min\ \{\operatorname{wind}_{\tau}(\nu)\mid\nu\ \text{eigenvalue},\ \nu\geq\delta\}$
$p^{\delta}(\gamma)$ $=\alpha^{\geq\delta}_{\tau}(\gamma)-\alpha^{<\delta}_{\tau}(\gamma)\in\{0,1\}\,.$
These numbers do not depend on $J$.
We define the constrained Conley–Zehnder index
(2.9) ${\rm CZ}^{\delta}_{\tau}(\gamma)=2\alpha^{<\delta}_{\tau}(\gamma)+p^{\delta}(\gamma).$
We note that ${\rm CZ}_{\tau}^{\delta}$ is a lower semi-continuous function of the asymptotic operator $L_{\gamma,J}$ with respect to the $C^{\infty}$ norm. We define the Conley-Zehnder index ${\rm CZ}_{\tau}(\gamma)={\rm CZ}_{\tau}^{0}(\gamma)$, i.e. by taking $\delta=0$ in (2.9).
If $\Omega\subset\mathbb{R}^{4}$ is a star-shaped domain and $Y=\partial\Omega$, then unless otherwise stated we choose $\tau$ to be a global trivialization of $\xi=\ker(\lambda_{0}|_{Y})$, and for any Reeb orbit $\gamma\in\mathcal{P}(Y)$, we define
${\rm CZ}(\gamma)={\rm CZ}_{\tau}^{0}(\gamma).$
2.3. Pseudoholomorphic curves in symplectizations
Let $J$ be a compatible complex structure on the symplectic vector bundle $(\xi,d\lambda)$. As in [93], consider an almost complex structure $\widetilde{J}$ defined on ×$\mathbb{R}\times Y$ by
(2.10) $\widetilde{J}\partial_{a}=R_{\lambda},\qquad\widetilde{J}|_{\xi}=J$
where $R_{\lambda}$ and $\xi$ are seen as $\mathbb{R}$-invariant objects in ×$\mathbb{R}\times Y$. Then $\widetilde{J}$ is $\mathbb{R}$-invariant. We say that an almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$ constructed this way is $\lambda$-compatible.
Consider a closed Riemann surface $(S,j)$, a finite set $\Gamma\subset S$ of “punctures”, and a holomorphic map
$\widetilde{u}=(a,u):(S\setminus\Gamma,j)\longrightarrow(\mathbb{R}\times Y,\widetilde{J})$
Define the Hofer energy
$E(\widetilde{u})=\sup_{\phi}\int_{S\setminus\Gamma}\widetilde{u}^{*}d(\phi\lambda).$
where the supremum is taken over the set of $\phi:\mathbb{R}\to[0,1]$ satisfying $\phi^{\prime}\geq 0$. If the Hofer energy $E(\widetilde{u})<\infty$, then we say that $\widetilde{u}$ is a finite energy curve.
A puncture $z\in\Gamma$ is positive or negative if $a(w)\to+\infty$ or $a(w)\to-\infty$ when $w\to z$, respectively. The puncture $z$ is removable if lim $\limsup|a(w)|<\infty$ when $w\to z$. It turns out that for a finite energy curve, every puncture is positive, negative or removable, and $\widetilde{u}$ can be smoothly extended across a removable puncture; see [93]. So we can assume without loss of generality that $\Gamma$ does not contain removable punctures, and we denote by $\Gamma_{+}$ and $\Gamma_{-}$ the sets of positive and negative punctures, respectively.
Let $z\in\Gamma$ and let $K\subset S$ be a conformal disk centered at $z$, meaning that there is a biholomorphism $\varphi:(K,j,z)\to(\mathbb{D},i,0)$. Then $K\setminus\{z\}$ admits positive holomorphic polar coordinates $(s,t)\in[0,+\infty)\times\mathbb{R}/\mathbb{Z}$ defined by $(s,t)\simeq\varphi^{-1}(e^{-2\pi(s+it)})$, and negative holomorphic polar coordinates $(s,t)\in(-\infty,0]\times\mathbb{R}/\mathbb{Z}$ defined by $(s,t)\simeq\varphi^{-1}(e^{2\pi(s+it)})$.
As at the beginning of §2, we say that $\lambda$ is nondegenerate up to action $C$ if all Reeb orbits of period less than $C$ are nondegenerate. If all Reeb orbits are nondegenerate, we say that $\lambda$ is nondegenerate.
Theorem 2.2 ([props1]).
Suppose that $\lambda$ is nondegenerate up to action $C$ and that $z$ is a nonremovable puncture of a finite energy curve $\widetilde{u}=(a,u)$ in (×$(\mathbb{R}\times Y,\widetilde{J})$ with Hofer energy $E(\widetilde{u})\leq C$. Let $(s,t)$ be positive holomorphic polar coordinates near $z$. There exist a Reeb orbit $\gamma$ with period $T$ and $d\in\mathbb{R}$ such that $u(s,t)\to\gamma(\epsilon Tt+d)$ in $C^{\infty}(\mathbb{R}/\mathbb{Z},Y)$ as $s\to+\infty$, where $\epsilon=\pm 1$ is the sign of the puncture.
The Reeb orbit $\gamma$ above is called the asymptotic limit of $\widetilde{u}$ at $z$.
Let $\gamma$ be a Reeb orbit of $\lambda$ with period $T$. We say that $\gamma$ is multiply covered if $\gamma$ is not simple, i.e. if $\gamma(T^{\prime})=\gamma(0)$ for some $T^{\prime}\in(0,T)$. In that case, the covering multiplicity of $\gamma$ is $k=T/T_{0}$ where $T_{0}$ is the minimal period.
Consider the space $\mathbb{R}/\mathbb{Z}\times\mathbb{C}$ equipped with coordinates $(\vartheta,z=x_{1}+ix_{2})$ and contact form $\beta_{0}=d\vartheta+x_{1}dx_{2}$.
Definition 2.3.
Let $\gamma$ be a Reeb orbit of $\lambda$ with period $T$ and convering multiplicity $k$. A Martinet tube for $\gamma$ is a smooth diffeomorphism $\Psi:\mathcal{N}\to\mathbb{R}/\mathbb{Z}\times\mathbb{D}$ defined on a smooth compact neighborhood $\mathcal{N}$ of $\gamma(\mathbb{R})$ such that:
  • $\Psi(\gamma(T\vartheta/k))=(\vartheta,0)$ for all $\vartheta\in\mathbb{R}/\mathbb{Z}$.
  • On $\mathcal{N}$ we have $\lambda=\Psi^{*}(g\beta_{0})$, where $g:\mathbb{R}/\mathbb{Z}\times\mathbb{D}\to(0,+\infty)$ is smooth and satisfies $g(\vartheta,0)=T/k$, $dg(\vartheta,0)=0$ for all $\vartheta\in\mathbb{R}/\mathbb{Z}$.
Theorem 2.4 ([props1, sie_CPAM]).
Suppose that $\lambda$ is nondegenerate up to action $C>0$ and that $z$ is a nonremovable puncture of sign $\epsilon=\pm 1$ of a finite energy curve $\widetilde{u}=(a,u)$ with Hofer energy $E(\widetilde{u})\leq C$. Let $(s,t)$ be positive or negative holomorphic polar coordinates at $z$ when $\epsilon=+1$ or $\epsilon=-1$, respectively. Let $\Psi:\mathcal{N}\to\mathbb{R}/\mathbb{Z}\times\mathbb{D}$ be a Martinet tube for the asymptotic limit $\gamma$ of $\widetilde{u}$ at $z$, and let $s_{0}\gg 1$ such that $|s|\geq s_{0}\Rightarrow u(s,t)\in\mathcal{N}$. Write $\Psi(u(s,t))=(\vartheta(s,t),z(s,t))$ for $|s|\geq s_{0}$. By applying a rotation we can assume $u(s,0)\to x(0)$ as $\epsilon s\to+\infty$.
If $z(s,t)$ does not vanish identically then the following holds. There exists $r>0$ and an eigenvalue $\nu$ of the asymptotic operator $L_{\gamma,J}$ satisfying $\epsilon\nu<0$, such that:
  • There exist $c,d\in\mathbb{R}$ and a lift $\tilde{\vartheta}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ of $\vartheta(s,t)$ such that
    $\lim_{\epsilon s\to+\infty}\sup_{t\in\mathbb{R}/\mathbb{Z}}e^{r\epsilon s}\left(|D^{\beta}[a(s,t)-Ts-c]|+|D^{\beta}[\tilde{\vartheta}(s,t)-kt-d]|\right)=0$
    holds for every partial derivative $D^{\beta}=\partial^{\beta_{1}}_{s}\partial^{\beta_{2}}_{t}$, where $k$ is the covering multiplicity of $\gamma$.
  • There exists an eigenvector of $\nu$, represented as a nowhere vanishing vector field $v(t)$ in the frame $\{\partial_{x_{1}},\partial_{x_{2}}\}$ along $\gamma$, such that
    $z(s,t)=e^{\nu s}(v(t)+R(s,t))$
    for some $R(s,t)$ satisfying $|D^{\beta}R(s,t)|\to 0$ in $C^{0}(\mathbb{R}/\mathbb{Z})$ as $\epsilon s\to+\infty$, for every partial derivative $D^{\beta}=\partial^{\beta_{1}}_{s}\partial^{\beta_{2}}_{t}$.
The eigenvalue $\nu$ provided by Theorem 2.4 is called the asymptotic eigenvalue of $\widetilde{u}$ at the puncture $z$.
Remark 2.5.
The alternative $z(s,t)\equiv 0$ can be expressed independently of coordinates as saying that the end of the domain of $\widetilde{u}$ corresponding to the puncture is mapped into the “trivial cylinder” ×$\mathbb{R}\times\gamma$ where $\gamma$ is the asymptotic limit. In this case we say that $\widetilde{u}$ has trivial asymptotic behaviour at the puncture. Otherwise, the asymptotic behaviour is said to be nontrivial at the puncture.
We now recall some topogical invariants introduced in [props2]. Let $\widetilde{u}=(a,u)$ be a finite energy curve in (×$(\mathbb{R}\times Y,\widetilde{J})$. Assume that $\lambda$ is nondegenerate up to action $E(\widetilde{u})$. Let
(2.11) $\pi_{\lambda}:TY\longrightarrow\xi$
denote the projection along the Reeb vector field $R_{\lambda}$. It can be shown using the Carleman similarity principle that if $\pi_{\lambda}\circ du$ does not vanish identically, then its zeros are isolated and count positively. Theorem 2.4 further implies that there are finitely many zeros in this case. So if $\widetilde{u}$ is nontrivial, then the algebraic count of zeros of $\pi_{\lambda}\circ du$ is a well-defined non-negative number, which following [props2] we denote by $\operatorname{wind}_{\pi}(\widetilde{u})$.
Fix a $d\lambda$-symplectic trivialization $\tau$ of $u^{*}\xi$. Let $z$ be a puncture of $\widetilde{u}$ with asymptotic limit $\gamma$ with period $T$. Let $\operatorname{wind}_{\infty}(\widetilde{u},z,\tau)=\operatorname{wind}_{\tau}(\nu)\in\mathbb{Z}$, where $\nu$ is the asymptotic eigenvalue of $\widetilde{u}$ at $z$. Finally, define
$\operatorname{wind}_{\infty}(\widetilde{u})=\sum_{z\in\Gamma_{+}}\operatorname{wind}_{\infty}(\widetilde{u},z,\tau)-\sum_{z\in\Gamma_{-}}\operatorname{wind}_{\infty}(\widetilde{u},z,\tau).$
Note that $\operatorname{wind}_{\infty}(\widetilde{u})$ does not depend on the choice of trivialization $\tau$ of $u^{*}\xi$. In fact, since $\pi_{\lambda}\circ du$ is a section of the complex line bundle $T^{1,0}S\otimes\xi$, it follows from degree theory that
(2.12) $\operatorname{wind}_{\pi}(\widetilde{u})=\operatorname{wind}_{\infty}(\widetilde{u})-\chi(S\setminus\Gamma).$
[…]
Definition 2.6 ([fast]).
The plane $\widetilde{u}$ is fast if $\operatorname{wind}_{\infty}(\widetilde{u})=1$.
Definition 2.7.
The covering multiplicity ${\rm cov}(\widetilde{u})$ of the plane $\widetilde{u}$ is the covering multiplicity of its asymptotic limit.
Lemma 2.8.
If $\widetilde{u}=(a,u)$ is a fast plane, then $\widetilde{u}$ is somewhere injective and the map $u:\mathbb{C}\to Y$ is an immersion which is transverse to $R_{\lambda}$.
Proof.
Since $\operatorname{wind}_{\pi}(\widetilde{u})\geq 0$, it follows from (2.12) that $\operatorname{wind}_{\pi}(\widetilde{u})=0$, and hence $u$ is an immersion which is transverse to $R_{\lambda}$. If $\widetilde{u}$ is not somewhere injective then it covers another plane via a polynomial map of degree $\geq 2$, and the covering map must have at least one critical point, contradicting the fact that $u$ is an immersion. ∎
2.4. Pseudoholomorphic curves in strong symplectic cobordisms
Let $(W,\omega)$ be a compact symplectic $4$-manifold for which there exists a Liouville vector field $V$ near $Y=\partial W$ that is transverse to $Y$. The $1$-form $\alpha=i_{V}\omega$ defined near $Y$ satisfies $d\alpha=\omega$ and restricts to a contact form on $Y$ which we denote by $\lambda$. We denote the associated contact structure by $\xi=\ker\lambda\subset TY$. The boundary splits as $Y=Y^{+}\sqcup Y^{-}$, where $V$ points out along $Y^{+}$ and points in along $Y^{-}$. The orientation of $Y$ as the boundary of $W$ agrees with the contact orientation on $Y^{+}$ and disagrees with it on $Y^{-}$. We call $(W,\omega)$ a strong symplectic cobordism that is convex at $(Y^{+},\lambda)$ and concave at $(Y^{-},\lambda)$. The local flow $\varphi^{t}_{V}$ of $V$ defines, for $\varepsilon>0$ small enough, diffeomorphisms
(2.13) $\Phi^{+}:(-\varepsilon,0]\times Y^{+}\longrightarrow\mathcal{U}^{+},\quad\Phi^{-}:[0,\varepsilon)\times Y^{-}\longrightarrow\mathcal{U}^{-},\quad\Phi^{\pm}(a,p)=\varphi^{a}_{V}(p)$
onto neighborhoods $\mathcal{U}^{\pm}$ of $Y^{\pm}$ in $W$. We have $(\Phi^{\pm})^{*}\alpha=e^{a}\lambda$, where here we write $\lambda$ for the pull-back of $\lambda$ under the projection ×$\mathbb{R}\times Y\to Y$. The symplectic form pulls back to
$\left(\Phi^{\pm}\right)^{*}\omega=d(e^{a}\lambda)=e^{a}(da\wedge\lambda+d\lambda).$
The symplectic completion of $(W,\omega)$ is the symplectic manifold $(\overline{W},\overline{\omega})$ defined by
(2.14) $\overline{W}=\left(\left(\left(-\varepsilon,+\infty\right)\times Y^{+}\right)\sqcup W\sqcup\left(\left(-\infty,\varepsilon\right)\times Y^{-}\right)\right)/\sim$
where points are identified according to
$\mathcal{U}^{+}\ni\Phi^{+}(a,p)\sim(a,p)\in(-\varepsilon,0]\times Y^{+},\quad\quad\mathcal{U}^{-}\ni\Phi^{-}(a,p)\sim(a,p)\in[0,\varepsilon)\times Y^{-}$
and the symplectic form $\overline{\omega}$ is defined by
$\overline{\omega}=\omega\ \text{on}\ W\qquad\qquad\overline{\omega}=d(e^{a}\lambda)\ \text{on}\ (-\varepsilon,+\infty)\times Y^{+}\ \text{and on}\ (-\infty,\varepsilon)\times Y^{-}.$
Fix a $\lambda$-compatible almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$. Consider an almost complex structure $\bar{J}$ on $\overline{W}$ that is $\omega$-compatible on $W$, and that agrees with $\widetilde{J}$ on $[0,+\infty)\times Y^{+}\cup(-\infty,0]\times Y^{-}$ as in (2.10). It follows that $\bar{J}$ is $\overline{\omega}$-compatible. As in [93, sft_comp] we consider a closed Riemann surface $(S,j)$, a finite set $\Gamma\subset S$ and a holomorphic map
$\widetilde{u}:(S\setminus\Gamma,j)\longrightarrow(\overline{W},\bar{J}).$
We assume the finite energy condition $0<E(\widetilde{u})<\infty$, where the energy $E(\widetilde{u})$ is defined by
$\begin{split}E(\widetilde{u})=&\int_{\widetilde{u}^{-1}(W)}\widetilde{u}^{*}\omega\\ &+\ \sup_{\phi}\left\{\int_{\widetilde{u}^{-1}([0,+\infty)\times Y^{+})}\widetilde{u}^{*}d(\phi\lambda)+\int_{\widetilde{u}^{-1}((-\infty,0]\times Y^{-})}\widetilde{u}^{*}d(\phi\lambda)\right\}.\end{split}$
Here the supremum is taken over the set of smooth $\phi:\mathbb{R}\to[0,1]$ satisfying $\phi^{\prime}\geq 0$. The finite energy condition implies that punctures are either removable or behave like the nonremovable punctures as explained in §2.3. More precisely, if $(s,t)$ are positive holomorphic polar coordinates at a positive puncture $z\in\Gamma$ then $\widetilde{u}(s,t)\in[0,+\infty)\times Y^{+}$ when $s\gg 1$, and if one writes $\widetilde{u}=(a,u)$ for $s\gg 1$ then $a(s,t)\to+\infty$ as $s\to+\infty$. There is a similar behaviour at negative punctures. Moreover, if $\lambda$ is nondegenerate up to action $E(\widetilde{u})$, then there is an asymptotic limit $P$ at a nonremovable puncture, and all the conclusions of Theorem 2.4 hold. In particular, we can talk about nontrivial versus trivial asymptotic behaviour and asymptotic eigenvalues, see Remark 2.5.
Now let $\widetilde{u}=(a,u):(\mathbb{C},i)\to(\overline{W},\bar{J})$ be a finite energy plane, where $\infty$ is a positive puncture [^3] Note that for a general $(W,\omega)$ the puncture of a plane might be negative. However it will always be a positive puncture when $\omega$ is exact.. Take $R>0$ large enough so that $u(\mathbb{C}\setminus B_{R}(0))\subset[0,+\infty)\times Y^{+}$. Then there is a splitting
(2.15) $\widetilde{u}^{*}T\overline{W}|_{\mathbb{C}\setminus B_{R}(0)}=(\widetilde{u}|_{\mathbb{C}\setminus B_{R}(0)})^{*}\left<\partial_{a},R_{\lambda}\right>\oplus(u|_{\mathbb{C}\setminus B_{R}(0)})^{*}\xi$
where $\partial_{a},R_{\lambda}\right>$ denotes the subbundle of $T([0,+\infty)\times Y^{+})$ spanned by $\{\partial_{a},R_{\lambda}\}$. There is a $d\lambda$-symplectic frame $\{e_{1},e_{2}\}$ of $(u|_{\mathbb{C}\setminus B_{R}(0)})^{*}\xi$ such that $\{\partial_{a},R_{\lambda},e_{1},e_{2}\}$ extends to a $\overline{\omega}$-symplectic frame of $\widetilde{u}^{*}T\overline{W}$. This frame is unique up to homotopy. If $\lambda$ nondegenerate up to action $E(\widetilde{u})$, then $\widetilde{u}$ has an asymptotic limit $\gamma=(x,T)\in\mathcal{P}(\lambda)$ in $Y^{+}$, and it follows from the asymptotic behaviour that, up to homotopy, $\{e_{1},e_{2}\}$ induces a $d\lambda$-symplectic trivialization $\tau_{\widetilde{u}}$ of $x(T\cdot)^{*}\xi$. Define
(2.16) ${\rm CZ}(\widetilde{u})={\rm CZ}_{\tau_{\widetilde{u}}}(\gamma),$
If the asymptotic behaviour of $\widetilde{u}$ is nontrivial, define
(2.17) $\operatorname{wind}_{\infty}(\widetilde{u})=\operatorname{wind}_{\tau_{\widetilde{u}}}(\nu)$
where $\nu$ is the asymptotic eigenvalue of $\widetilde{u}$ at the puncture $\infty$. If the asymptotic behaviour is trivial then we define
(2.18) $\operatorname{wind}_{\infty}(\widetilde{u})=-\infty.$
Under the above nondegeneracy assumption, we define a plane with a positive puncture to be fast if $\operatorname{wind}_{\infty}(\widetilde{u})\leq 1$. Unlike the case of symplectizations, here a fast plane need not be immersed or somewhere injective; the proof of Lemma 2.8 does not carry over because $\operatorname{wind}_{\pi}$ does not make sense in this context. As before, we define ${\rm cov}(\widetilde{u})$ to be the covering multiplicity of the asymptotic limit. If ${\rm cov}(\widetilde{u})=1$ then $\widetilde{u}$ is somewhere injective.
2.5. The main compactness argument
We now prove Lemma 2.1. We proceed in five steps.
Step 1
We begin with some geometric setup.
Consider the compact manifold
$W=\Omega_{2}\setminus\varphi(\Omega_{1}\setminus\partial\Omega_{1}),$
with the symplectic form $\omega$ given by the restriction of the standard symplectic form $\omega_{0}$ on $\Omega_{2}$. We have
$\partial W=Y^{+}-Y^{-},$
where $Y^{+}=\partial\Omega_{2}$ and $Y^{-}=\varphi(\partial\Omega_{1})$. The radial vector field on $\Omega_{2}$ is Liouville, transverse and outward pointing at $Y^{+}$. The pushforward of the radial vector field on $\Omega_{1}$ by $\varphi$ extends to a Liouville vector field defined on a neighbourhood of $Y^{-}$ in $W$, where it is transverse to $Y^{-}$ and inward pointing. Hence we can define the symplectic completion $(\overline{W},\overline{\omega})$ of $(W,\omega)$ as explained in subsection 2.4. As usual, let $\lambda$ denote the contact form on $Y^{+}\cup Y^{-}$. By assumption, $\lambda$ is, up to action $C$, dynamically convex and nondegenerate. Write $\xi=\ker\lambda$. Fix a global $d\lambda$-symplectic trivialization $\tau$ of $\xi$.
Choose any complex structure $J$ on $\xi$ which is $d\lambda$-compatible, and let $\widetilde{J}$ be the corresponding $\lambda$-compatible almost complex structure on ×$\mathbb{R}\times Y$ as in equation (2.10). Let $\bar{J}$ be an almost complex structure on $\overline{W}$ which is $\omega$-compatible on $W$ and agrees with $\widetilde{J}$ on $(-\infty,0]\times Y^{-}\cup[0,+\infty)\times Y^{+}$. In particular, $\bar{J}$ is $\overline{\omega}$-compatible. We need to choose the extension $\bar{J}$ generically, in a way to be specified in Step 4 below.
Step 2
Let $\gamma_{+}\in\mathcal{P}(Y^{+})$ be a simple Reeb orbit with period $T\leq C$. Let $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ denote the moduli space of equivalence classes of fast and embedded finite energy planes $\widetilde{u}:(\mathbb{C},i)\to(\overline{W},\bar{J})$ asymptotic to $\gamma_{+}$. Here two planes $\widetilde{u}_{0},\widetilde{u}_{1}$ are declared equivalent if there exist $(A,B)\in\mathbb{C}^{*}\times\mathbb{C}$ such that $\widetilde{u}_{1}(Az+B)=\widetilde{u}_{0}(z)$ for all $z\in\mathbb{C}$. We now show that $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ is naturally endowed with the structure of a two-dimensional smooth manifold.
Fix $\delta<0$ in the interval between eigenvalues of the asymptotic operator $L_{\gamma_{+},\widetilde{J}}$ satisfying $\operatorname{wind}_{\tau}=1$ and $\operatorname{wind}_{\tau}=2$. This is possible since it follows from the definition of fast plane that ${\rm CZ}^{0}_{\tau}(\gamma_{+})\geq 3$. Note that $\alpha^{<\delta}_{\tau}(\gamma_{+})=1$ and ${\rm CZ}^{\delta}_{\tau}(\gamma_{+})=3$. Following [props3], one can build a Fredholm theory with weight $\delta$ based on sections of the normal bundle. In more detail, let $\widetilde{u}=(a,u)$ represent an element of $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$. Using the asymptotic behaviour from Theorem 2.4 and the fact that $\widetilde{u}$ is assumed to be an embedding (an immersion would suffice), one can consider a $\bar{J}$-invariant normal bundle $N$ of $\widetilde{u}$ that agrees with $u^{*}\xi$ near the puncture $\infty$. Then one looks at sections of the normal bundle that decay faster then $e^{\delta s}$ near $\infty$, where a neighbourhood of $\infty$ is parametrised by $(s,t)\simeq e^{2\pi(s+it)}$, $s\gg 1$. Here one can use Sobolev or Hölder spaces. The linearization of the Cauchy-Riemann equations at $\widetilde{u}$ determines a Fredholm operator $D_{\widetilde{u}}$ on this space of sections with index
${\rm ind}_{\delta}(\widetilde{u})={\rm CZ}^{\delta}_{\tau}(\gamma_{+})-1=3-1=2.$
We claim that automatic transversality holds, i.e. this Fredholm operator is always surjective, without any genericity assumption on the almost complex structure. To see this, note that a symplectic trivialization of the normal bundle can always be homotoped to a trivialization that extends over a $d\lambda$-symplectic trivialization $\tau_{N}$ of $x(T\cdot)^{*}\xi$. The decomposition of complex vector bundles $\widetilde{u}^{*}TW=T\mathbb{C}\oplus N$ implies that a section of $x(T\cdot)^{*}\xi$ with winding number equal to 00 with respect to $\tau_{N}$ will have winding number equal to $+1$ with respect to $\tau$. In particular, $\alpha^{<\delta}_{\tau_{N}}(\gamma_{+})=\alpha^{<\delta}_{\tau}(\gamma_{+})-1=1-1=0$. A section $\zeta$ in the kernel of $D_{\widetilde{u}}$ satisfies a Cauchy-Riemann type equation. Hence either $\zeta\equiv 0$, or its zeros are isolated and count positively to the total algebraic count of zeros. Moreover, if $\zeta\not\equiv 0$ then its asymptotic behaviour at $\infty$ is governed by an eigenvector $e(t)$ of the asymptotic operator $L_{\gamma_{+},\widetilde{J}}$ associated to an eigenvalue $\nu<0$ in a similar manner to what is described by Theorem 2.4: $\zeta(s,t)\sim e^{\nu s}(e(t)+\varepsilon(s,t))$, $\sup_{t}|\varepsilon(s,t)|\to 0$ as $s\to+\infty$. One can justify such asymptotic behaviour using [sie_CPAM, Theorem A.1]. Hence $\zeta$ has only finitely many zeros, and by the argument principle the total count of zeros of $\zeta$ is nonnegative and equal to $\operatorname{wind}_{\tau_{N}}(\nu)$. The exponential decay $e^{\delta s}|\zeta(s,t)|\to 0$ forces $\nu<\delta$. Thus $\operatorname{wind}_{\tau_{N}}(\nu)\leq\alpha^{<\delta}_{\tau_{N}}(\gamma_{+})=0$, and we conclude that $\zeta$ never vanishes if it does not vanish identically. If $D_{\widetilde{u}}$ is not surjective then its kernel has dimension $\geq 3$ since the index is $2$, so we would find three linearly independent sections in the kernel, but this is impossible: a linear combination of them would vanish at some point because the normal bundle has rank two. This proves the desired surjectivity.
The automatic transversality proved above endows $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ with the structure of a two-dimensional manifold. Note that since the Fredholm theory is modelled on a space of sections with exponential decay faster than $\delta$, the asymptotic eigenvalue of the obtained nearby planes is always $<\delta$, implying that $\operatorname{wind}_{\infty}\leq\alpha^{<\delta}_{\tau}=1$ for the nearby planes.
In preparation for compactness arguments below, we note that one can use results on cylinders of small area such as those proven in [small_area], together with the assumed nondegeneracy of $\gamma_{+}$, to conclude that the topology $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ inherits from the functional analytic set-up agrees with the topology induced by $C^{\infty}_{\rm loc}$-convergence. This is the property of “completeness” in [props3].
Step 3
Suppose now that $\widetilde{u}$ is a embedded fast plane in $\overline{W}$ asymptotic to $\gamma_{+}$. Let $\mathcal{Y}\subset\mathcal{M}^{\rm fast}(\gamma,\bar{J})$ denote the connected component of $[\widetilde{u}]$. For each $L\geq 0$ consider the compact subset of $\overline{W}$ defined by
$E_{L}=\left([-L,0]\times Y^{-}\right)\sqcup W\sqcup\left([0,L]\times Y^{+}\right).$
Define
$\mathcal{Y}_{L}=\{[\widetilde{v}]\in\mathcal{Y}\mid\widetilde{v}(\mathbb{C})\cap E_{L}\neq\emptyset,\ \widetilde{v}(\mathbb{C})\cap\left((-\infty,L)\times Y^{-}\right)=\emptyset\}.$
The goal of this step is to show that $\mathcal{Y}_{L}$ is compact.
Let $[\widetilde{v}_{n}]\in\mathcal{Y}_{L}$ be an arbitrary sequence. By the SFT compactness theorem [sft_comp] we find a subsequence, still denoted by $\widetilde{v}_{n}$, that SFT converges to a holomorphic building ${\bf u}$. The bottom level of the building is a (possibly nodal) curve in $(\overline{W},\bar{J})$, but in principle there might be more levels in (×$(\mathbb{R}\times Y^{+},\widetilde{J})$. Any level in ×$\mathbb{R}\times Y^{+}$ must have at least one irreducible component which is not a trivial cylinder ×$\mathbb{R}\times\gamma$ where $\gamma$ is a Reeb orbit.
We first prove that ${\bf u}$ has only the bottom level. Suppose to get a contradiction that ${\bf u}$ has levels in ×$\mathbb{R}\times Y^{+}$. Then the top level is a finite energy punctured sphere
$\widetilde{u}_{+}=(a_{+},u_{+}):(\mathbb{C}\setminus\Gamma,i)\longrightarrow(\mathbb{R}\times Y^{+},\widetilde{J})$
[…]
We now claim that
(2.19) $\operatorname{wind}_{\infty}(\widetilde{u}_{+},\infty,\tau)\leq 1.$
To prove (2.19), SFT compactness provides constants $A_{n}\in\mathbb{C}^{*}$, $B_{n}\in\mathbb{C}$ such that the sequence of maps $\widetilde{w}_{n}(z)=\widetilde{v}_{n}(A_{n}z+B_{n})$ has the following properties:
  • (i)
    For every open neighborhood $U$ of $\Gamma$ in $\mathbb{C}$ and every $A\geq 0$ there exists $n_{U,A}$ such that if $n\geq n_{U,A}$ then $\widetilde{w}_{n}(\mathbb{C}\setminus U)\subset[A,+\infty)\times V^{+}$.
  • (ii)
    By (i) we can write $\widetilde{w}_{n}=(d_{n},w_{n})$ on arbitrary compact subsets of $\mathbb{C}\setminus\Gamma$ provided $n$ large enough. Then $w_{n}\to u_{+}$ in $C^{\infty}_{\rm loc}(\mathbb{C}\setminus\Gamma)$.
Note that there is a uniformity in neighbourhoods of $\infty$ encoded in property (i). Write $(s,t)$ instead of $e^{2\pi(s+it)}$. Theorem 2.4 provides $s_{0}>0$ such that $\pi_{\lambda}(\partial_{s}u_{+})$ does not vanish on $[s_{0},+\infty)\times\mathbb{R}/\mathbb{Z}$. In particular the winding number $\operatorname{wind}_{\tau}(\pi_{\lambda}(\partial_{s}u_{+}(s_{0},\cdot)))$ of $t\mapsto\pi_{\lambda}(\partial_{s}u_{+}(s_{0},t))$ in the global frame $\tau$ is equal to $\operatorname{wind}_{\infty}(\widetilde{u},\infty,\tau)$. By (i) we can assume, up to making $s_{0}$ larger, that $\pi_{\lambda}(\partial_{s}w_{n})$ is defined on the set $[s_{0},+\infty)\times\mathbb{R}/\mathbb{Z}$. Use $\tau$ to represent $\pi_{\lambda}(\partial_{s}w_{n})$ as a $\mathbb{C}$-valued function $\zeta_{n}(s,t)$. This function satisfies a Cauchy-Riemann type equation, hence its zeros count positively towards the total algebraic count of zeros. Since $\pi_{\lambda}(\partial_{s}w_{n}(s_{0},\cdot))\to\pi_{\lambda}(\partial_{s}u_{+}(s_{0},\cdot))$ in $C^{\infty}(\mathbb{R}/\mathbb{Z})$ and the latter does not vanish, we get that $\zeta_{n}(s_{0},\cdot)$ does not vanish and $\operatorname{wind}(\zeta_{n}(s_{0},\cdot))=\operatorname{wind}_{\infty}(\widetilde{u}_{+},\infty,\tau)$ for all $n$ large enough. The argument principle now implies that
$1\geq\operatorname{wind}_{\infty}(\widetilde{v}_{n})=\lim_{s\to+\infty}\operatorname{wind}(\zeta_{n}(s,\cdot))\geq\operatorname{wind}(\zeta_{n}(s_{0},\cdot))=\operatorname{wind}_{\infty}(\widetilde{u}_{+},\infty,\tau)$
for all $n\gg 1$, as desired.
With (2.19) established, we now invoke the dynamical convexity of $\lambda$ up to action $C$. Namely, at each negative puncture $z\in\Gamma$ we have $\operatorname{wind}_{\infty}(\widetilde{u},z,\tau)\geq 2$ since the corresponding asymptotic limits satisfy ${\rm CZ}^{0}_{\tau}\geq 3$. Hence
$0\leq\operatorname{wind}_{\pi}(\widetilde{u}_{+})=\operatorname{wind}_{\infty}(\widetilde{u}_{+})-1+\#\Gamma\leq 1-2\#\Gamma-1+\#\Gamma=-\#\Gamma,$
so $\Gamma=\emptyset$. This contradiction concludes the proof of the claim that ${\bf u}$ has only one level.
The bottom level of ${\bf u\/}$ cannot contain any closed irreducible components, because the symplectic form $\overline{\omega}$ on $\overline{W}$ is exact. Thus we have shown that a sequence $\widetilde{v}_{n}$ representing a sequence in $\mathcal{Y}_{L}$ has $C^{\infty}_{\rm loc}$-convergent subsequence, up to reparametrisations, to a finite-energy plane $\widetilde{v}$ asymptotic to $\gamma_{+}$.
Denoting again by $\widetilde{v}_{n}$ the convergent subsequence, the SFT compactness provides information analogous to (i) and (ii):
  • (iii)
    There exists some open bounded set $U\subset\mathbb{C}$ and some $N$ such that if $n\geq N$ then $\widetilde{v}_{n}(\mathbb{C}\setminus U)\subset[0,+\infty)\times Y^{+}$.
Now the exact same argument as above will show that $\operatorname{wind}_{\infty}(\widetilde{v})\leq 1$, i.e. $\widetilde{v}$ is fast. Since $\gamma$ is simply covered we know that $\widetilde{v}$ is somewhere injective. Results from [MW] now imply that $\widetilde{v}$ has to be an embedding, since critical points or self-intersections of $\widetilde{v}$ would force that $\widetilde{v}_{n}$ is not an embedding for $n$ large enough, a contradiction. This concludes the proof that $\mathcal{Y}_{L}$ is compact.
Step 4
We now produce the Reeb orbit $\gamma$ claimed by the Lemma 2.1, which will be denoted by $\gamma_{-}$ below.
The Fredholm theory from Step 2 provides families of planes in $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ by exponentiating nonvanishing sections of the normal bundle. Hence the set of points in $\overline{W}$ contained in the image of some plane in $\mathcal{Y}$ is open. This fact and the compactness of $\mathcal{Y}_{L}$ proved above together show that we can find a sequence of planes $[\widetilde{v}_{n}]\in\mathcal{Y}$ satisfying
(2.20) $\forall L\geq 0\ \exists n_{L}\ :\ n\geq n_{L}\Longrightarrow\widetilde{v}_{n}(\mathbb{C})\cap\left((-\infty,-L]\times Y^{-}\right)\neq\emptyset.$
By SFT compactness again, we can pass to a subsequence so that $[\widetilde{v}_{n}]$ converges to a building ${\bf u}$. Unlike in Step 3, ${\bf u}$ now has at least one lower level in (×$(\mathbb{R}\times Y^{-},\widetilde{J})$, in addition to a unique level in $(\overline{W},\bar{J})$, and possibly some upper levels in (×$(\mathbb{R}\times Y^{+},\lambda)$.
The same argument as in Step 3, using (i) and (ii), shows that there are no upper levels in (×$(\mathbb{R}\times Y^{+},\lambda)$. It follows that the level in $(\overline{W},\bar{J})$ consists precisely of a finite energy sphere $\widetilde{u}_{0}:(\mathbb{C}\setminus\Gamma_{0},i)\to(\overline{W},\bar{J})$ with precisely one positive puncture (at $\infty$) where the asymptotic limit is $\gamma_{+}$. If the asymptotic behaviour at the positive puncture is trivial then $\operatorname{wind}_{\infty}(\widetilde{u}_{0},\tau,\infty)=-\infty$. If the the asymptotic behaviour at $\infty$ is nontrivial trivial then, up to reparametrisation, the SFT compactness theorem provides property (iii), and one can argue similarly to Step 3 to prove that $\operatorname{wind}_{\infty}(\widetilde{u}_{0},\tau,\infty)\leq 1$.
We claim now that $\widetilde{u}_{0}$ is an embedded cylinder, asymptotic at its negative puncture to a Reeb orbit $\gamma_{-}$ satisfying ${\rm CZ}(\gamma_{-})=3$. To see this, first note that $\widetilde{u}_{0}$ is somewhere injectiv,e because the asymptotic limit at the unique positive puncture $\infty$ is the simple Reeb orbit $\gamma_{+}$. Since it is a limit of embeddings, the results from [MW] imply that $\widetilde{u}_{0}$ is also an embedding. Now one builds a weighted Fredholm theory with sections of the normal bundle for embedded curves in $(\overline{W},\bar{J})$ with the same asymptotic behaviour as $\widetilde{u}_{0}$. There is only the nontrivial weight $\delta<0$ at the positive where the asymptotic limit is $\gamma_{+}$. The weighted Fredholm index is
${\rm CZ}^{\delta}_{\tau}(\gamma_{+})-\sum_{z\in\Gamma_{0}}{\rm CZ}^{0}_{\tau}(\gamma_{z})-1+\#\Gamma_{0}$
Here $\gamma_{z}$ denotes the asymptotic limit of $\widetilde{u}_{0}$ at the negative punctures $z\in\Gamma_{0}$. Standard arguments imply that if $\bar{J}$ is chosen in a residual subset of almost complex structures, then this the weighted Fredholm index is nonnegative. Assume that $\bar{J}$ has been chosen this way in Step 1. We know that ${\rm CZ}^{\delta}_{\tau}(\gamma)=3$ and ${\rm CZ}^{0}_{\tau}(\gamma_{z})\geq 3$ by dynamical convexity. Plugging this in to the above formula for the weighted Fredholm index, we obtain
(2.21) $0\leq 3-3\#\Gamma_{0}-1+\#\Gamma_{0}=2(1-\#\Gamma_{0}).$
It follows that #$\#\Gamma_{0}=1$ and the above inequalities are equalities. Thus $\widetilde{u}_{0}$ is an embedded cylinder asymptotic at its negative puncture to a Reeb orbit $\gamma_{-}$ satisfying ${\rm CZ}(\gamma_{-})=3$.
Step 5
We claim now that $\gamma_{-}$ fulfills the requirements of Lemma 2.1. We have
$\mathcal{A}(\gamma_{-})<\mathcal{A}(\gamma_{+})\leq\mathcal{A}_{\operatorname{Hopf}}(\Omega_{2})$
by Stokes theorem since $(W,\omega)$ is exact. So we just need to show that $\gamma_{-}$ is a Hopf orbit.
[…]
(2.22) $0\leq\operatorname{wind}_{\pi}(\widetilde{u}_{-})=\operatorname{wind}_{\infty}(\widetilde{u}_{-})-1+\#\Gamma_{-}\leq 1-2\#\Gamma_{-}-1+\#\Gamma_{-}=-\#\Gamma_{-}$
implying that $\Gamma_{-}=\emptyset$ and $\widetilde{u}_{-}$ is a plane. It is automatically fast: $\operatorname{wind}_{\infty}(\widetilde{u}_{-})=\operatorname{wind}_{\infty}(\widetilde{u}_{-},\infty,\tau)\leq 1$. It remains to show that $\widetilde{u}_{-}$ is embedded, but this follows from the fact that it is a somewhere injective limit of embeddings.
In [fast, hryn_jsg] it is proved444 In the light of results from [sief_int] this can now be explained in more modern language. If the embedded fast plane $\widetilde{u}=(a,u)$ in (×$(\mathbb{R}\times\partial W,\widetilde{J})$ has as asymptotic limit a simply covered periodic orbit $\hat{P}$ then one can define a weighted self-intersection number $\widetilde{u}*_{\delta}\widetilde{u}={\rm int}(\widetilde{u},\widetilde{u}_{\tau})+\alpha^{<\delta}_{\tau}(\hat{P})$. Here $\tau$ is any $d\lambda$-symplectic trivialization of $\xi$ along $\hat{P}$, and $\widetilde{u}_{\tau}=(a,u_{\tau})$ where $u_{\tau}$ is the map $u$ pushed in the direction of $\tau$ near the puncture. Arguing as in [sief_int], it follows that $\widetilde{u}*_{\delta}\widetilde{u}\geq 0$ and that if $\widetilde{u}*_{\delta}\widetilde{u}=0$ then $u$ is an embedding into $\partial W\setminus\hat{P}$. Plugging $\tau=\tau_{N}$, where $\tau_{N}$ is the trivialization that extends over the normal bundle, we get ${\rm int}(\widetilde{u},\widetilde{u}_{\tau_{N}})=0$ and $\alpha^{<\delta}_{\tau}(\hat{P})=\alpha^{<\delta}_{\tau}(\hat{P})-1=1-1=0$. Hence $\widetilde{u}*_{\delta}\widetilde{u}=0$ and $\hat{P}$ is unknotted. Since $u$ is transverse to the Reeb vector field, the self-linking number of $\hat{P}$ is $-1$. that a simple Reeb orbit that is the asymptotic limit of an embedded fast finite energy plane is unknotted and has self-linking number $-1$. Thus $\gamma_{-}$ is a Hopf orbit, and this completes the proof of Lemma 2.1.
3. The first ECH capacity from a Hopf orbit
We now recall the definition of $c_{1}^{\rm ECH}$ and prove Theorem 1.14.
3.1. Embedded contact homology
To begin, let $Y$ be a closed three-manifold and let $\lambda$ be a nondegenerate contact form on $Y$. We now review the definition of embedded contact homology ${\rm ECH}(Y,\lambda)$. For simplicity we further assume that $Y$ is a homology sphere, which is sufficient for our purposes; for more details and the case of general three-manifolds see e.g. [ind_rev, 26].
Definition 3.1.
An ECH generator is a finite set of pairs $\alpha=\{(\alpha_{i},m_{i})\}$ where the $\alpha_{i}$ are distinct simple Reeb orbits, the $m_{i}$ are positive integers, and $m_{i}=1$ whenever $\alpha_{i}$ is hyperbolic [^5] A nondegenerate Reeb orbit is hyperbolic if the linearized return map (2.1) has real eigenvalues.. We sometimes denote an ECH generator as a product $\alpha=\prod_{i}\alpha_{i}^{m_{i}}$.
Definition 3.2.
If $\alpha=\{(\alpha_{i},m_{i})\}$ is an ECH generator, its symplectic action is
$\mathcal{A}(\alpha)=\sum_{i}m_{i}\mathcal{A}(\alpha_{i}).$
Definition 3.3.
If $\alpha=\alpha_{1}^{m_{1}}\cdots\alpha_{k}^{m_{k}}$ is an ECH generator, its ECH index is the integer $I(\alpha)$ defined by [^6] To compare with the notation in other papers: $I(\alpha)$ here corresponds to $I(\alpha,\emptyset)$ in [ind_rev, Def. 2.15], and to $I(\alpha,\emptyset,Z)$ in [26, Eq. (3.4)], where $Z$ is the unique relative homology class of surface in $Y$ with boundary $\alpha$.
(3.1) $I(\alpha)=\sum_{i}\left(m_{i}c_{\tau}(\alpha_{i})+m_{i}^{2}Q_{\tau}(\alpha_{i})\right)+\sum_{i\neq j}m_{i}m_{j}\operatorname{link}(\alpha_{i},\alpha_{j})+\sum_{i}\sum_{l=1}^{m_{i}}{\rm CZ}_{\tau}(\alpha_{i}^{l}).$
Here $\tau$ is any trivialization of $\xi$ over the Reeb orbits $\alpha_{i}$; the ECH index $I(\alpha)$ does not depend on this choice, although the terms on the right hand side of (3.1) do. If $\gamma$ is a simple Reeb orbit, then $c_{\tau}(\gamma)$ denotes the relative first Chern class of $\xi$ with respect to $\tau$ over a surface bounded by $\gamma$, and $Q_{\tau}(\gamma)$ denotes the relative self-intersection number with respect to $\tau$ of a surface bounded by $\gamma$; see [ind_rev, 26] for details. Also, $\operatorname{link}$ denotes the linking number, and $\gamma^{l}$ denotes the Reeb orbit that covers $\gamma$ with multiplicity $l$.
Remark 3.4.
In our situation it is useful to rewrite the index formula (3.1) as follows. First, it follows from the definitions that if $\gamma$ is a simple Reeb orbit, then for any trivialization $\tau$ of $\xi$ over $\gamma$, the self-linking number of $\gamma$ satisfies
(3.2) $\operatorname{sl}(\gamma)=Q_{\tau}(\gamma)-c_{\tau}(\gamma).$
Second, if we choose $\tau$ to be a global trivialization of $\xi$ (which we can do, uniquely up to homotopy, since $Y$ is a homology sphere), then $c_{\tau}(\gamma)=0$ for every Reeb orbit $\gamma$. Consequently, we can rewrite equation (3.1) as
(3.3) $I(\alpha)=\sum_{i}m_{i}^{2}{\rm sl}(\alpha_{i})+\sum_{i\neq j}m_{i}m_{j}\operatorname{link}(\alpha_{i},\alpha_{j})+\sum_{i}\sum_{l=1}^{m_{k}}{\rm CZ}(\alpha_{i}^{l}).$
Definition 3.5.
Define $\operatorname{ECC}(Y,\lambda)$ to be the vector space [^7] It is also possible to define ECH with $\mathbb{Z}$ coefficients, but we do not need this. over $\mathbb{Z}/2$ freely generated by the ECH generators, with $\mathbb{Z}$ grading defined by (3.3).
Next, choose a generic $d\lambda$-compatible almost complex structure on $\xi$, and let $\widetilde{J}$ denote the corresponding almost complex structure on ×$\mathbb{R}\times Y$ as in §2.3. If $\alpha$ and $\beta$ are ECH generators, define $\mathcal{M}^{J}(\alpha,\beta)$ to be the moduli space of $\widetilde{J}$-holomorphic currents [^8] A “holomorphic current” is a finite linear combination of somewhere injective holomorphic curves with positive integer multiplicities; see [26, §3.1] for details. in ×$\mathbb{R}\times Y$ asymptotic to $\alpha$ as the $\mathbb{R}$ coordinate goes to +$\infty$+ ∞ and asymptotic to $\beta$ as the $\mathbb{R}$ coordinate goes to $\infty$- ∞. If $J$ is generic, then when $I(\alpha)-I(\beta)=1$, the moduli space $\mathcal{M}^{J}(\alpha,\beta)/\mathbb{R}$ is finite, where $\mathbb{R}$ acts by translation of the $\mathbb{R}$ coordinate on ×$\mathbb{R}\times Y$; see [26, §5.3].
Definition 3.6.
If $J$ is generic as above, we define a differential
$\partial_{J}:ECC_{*}(Y,\lambda)\longrightarrow ECC_{*-1}(Y,\lambda)$
as follows. If $\alpha$ is an ECH generator, then
$\partial_{J}\alpha=\sum_{I(\alpha)-I(\beta)=1}\#\frac{\mathcal{M}^{J}(\alpha,\beta)}{\mathbb{R}}\beta$
where $\beta$ is an ECH generator and ##\## denotes the mod 2 count.
It is shown in [HT1, Thm. 7.20] that $\partial_{J}^{2}=0$. We denote the homology of the chain complex $(ECC_{*}(Y,\lambda),\partial_{J})$ by $ECH_{*}(Y,\lambda,J)$. It was shown by Taubes [taubes_iso] that $ECH_{*}(Y,\lambda,J)$ is canonically isomorphic to a version of Seiberg-Witten Floer cohomology as defined by Kronheimer-Mrowka [32]. As explained in [31, Thm. 1.3], this implies that $ECH_{*}(Y,\lambda,J)$ in fact depends only on the pair $(Y,\xi)$, so we can denote it by $ECH_{*}(Y,\xi)$.
The case of interest for us is where $Y=S^{3}$ and $\xi$ is the “standard” contact structure (the unique tight contact structure) $\xi_{0}$ arising from realizing $S^{3}$ as a star-shaped hypersurface in $\mathbb{R}^{4}$ with the restriction of the standard Liouville form. In this case we have
(3.4) $\operatorname{ECH}_{*}(S^{3},\xi_{0})=\left\{\begin{array}[]{cl}\mathbb{Z}/2,&*=0,2,4,\ldots\\ 0,&\mbox{else}.\end{array}\right.$
See e.g. the calculation in [26, §3.7].
There is some additional structure on ECH which we will need. First of all, the empty set is a valid ECH generator, with grading $I(\emptyset)=0$. For any $J$ we have $\partial_{J}\emptyset=0$, because the ECH differential decreases action, as explained in §3.2 below. It follows from the properties of ECH cobordism maps in [31, Thm. 1.9] that the homology class $[\emptyset]\in\operatorname{ECH}_{0}(Y,\lambda,J)$ gives a well-defined invariant
$c(\xi)\in\operatorname{ECH}_{0}(Y,\xi).$
In the example (3.4), $c(\xi_{0})$ is the generator of $\operatorname{ECH}_{0}(S^{3},\xi_{0})$.
There is also a “$U$ map”
$U:\operatorname{ECH}_{*}(Y,\xi)\longrightarrow\operatorname{ECH}_{*-2}(Y,\xi).$
This is induced by a chain map
$U_{J,y}:\operatorname{ECC}_{*}(Y,\lambda)\longrightarrow\operatorname{ECC}_{*-2}(Y,\lambda)$
which counts $\widetilde{J}$-holomorphic currrents passing through $(0,y)$ where $y\in Y$ is a point not on any Reeb orbit. See [26, §3.8] for details about why the $U$ map is well defined. In the example (3.4), the $U$ map on $\operatorname{ECH}_{*}(S^{3},\xi_{0})$ is an isomorphism whenever *0$\neq$ 0* ≠ 0.
3.2. ECH capacities
Choose a generic almost complex structure $J$ as needed to define the ECH differential $\partial_{J}$ on $\operatorname{ECC}_{*}(Y,\lambda)$. It follows from the definition of the almost complex structure $\widetilde{J}$ that if $\beta$ is another ECH generator and if $\mathcal{M}^{J}(\alpha,\beta)\neq\emptyset$, then $\mathcal{A}(\alpha)\geq\mathcal{A}(\beta)$, with equality only if $\alpha=\beta$. In particular, if $\beta$ appears in $\partial_{J}\alpha$, then
(3.5) $\mathcal{A}(\alpha)>\mathcal{A}(\beta).$
It follows that for each $L\in\mathbb{R}$, we have a subcomplex $\operatorname{ECC}_{*}^{L}(Y,\lambda)$ of $\operatorname{ECC}_{*}(Y,\lambda)$, spanned by ECH generators with symplectic action less than $L$.
Definition 3.7.
The filtered ECH, denoted by $\operatorname{ECH}^{L}(Y,\lambda)$, is the homology of the chain complex $(\operatorname{ECC}_{*}^{L}(Y,\lambda),\partial_{J})$.
It is shown in [31, Thm. 1.3] that the filtered ECH depends only on $Y,\lambda,L$ and not on the choice of $J$. In addition, inclusion of chain complexes inducs a well-defined chain map
$\imath_{*}^{L}:\operatorname{ECH}_{*}^{L}(Y,\lambda)\longrightarrow\operatorname{ECH}_{*}(Y,\lambda).$
Definition 3.8.
Let $X\subset\mathbb{R}^{4}$ be a star-shaped domain, let $Y=\partial X$, let $\lambda$ be the contact form on $Y$ given by the restriction of the standard Liouville form, and suppose that $\lambda$ is nondegenerate. If $k$ is a nonnegative integer, define the $k^{th}$ ECH capacity of $X$ by
$c_{k}(X)=\inf\{L>0\mid 0\neq\imath_{2k}^{L}:\operatorname{ECH}_{2k}^{L}(Y,\lambda)\to\operatorname{ECH}_{2k}(Y,\lambda)\}.$
Equivalently, if $J$ is an almost complex structure as needed to define the ECH generator $\partial_{J}$, then $c_{k}(X)$ is the minimum $L$ such that the generator of $ECH_{2k}(Y,\lambda,J)$ can be represented by a sum of ECH generators each having action at most $L$.
The capacity $c_{k}$ has a unique $C^{0}$-continuous extension to the space of all star-shaped domains in $\mathbb{R}^{4}$; see [hut_qech, §4] for explanation and for the definition of ECH capacities of more general symplectic manifolds.
3.3. Proof of Theorem 1.14
Let $X\subset\mathbb{R}^{4}$ be a dynamically convex domain. Let $Y=\partial X$ and let $\lambda$ denote the induced contact form on $Y$. By continuity of capacities, we can assume that $\lambda$ is nondegenerate. We now prove that $c_{\operatorname{Hopf}}(X)=c_{1}^{\operatorname{ECH}}(X)$ in three steps.
Step 1
We first prove that $c_{\rm Hopf}(X)\geq c_{1}^{{\rm ECH}}(X)$. One can quickly deduce this step from the result of Edtmair [gss_edtmair] that $c_{Z}(X)\leq\mathcal{A}_{\operatorname{Hopf}}(X)$, but we will give an independent proof.
By Theorem 1.12, there exists a Hopf orbit $\gamma\in\mathcal{P}_{\operatorname{Hopf}}(\partial X)$ satisfying $c_{\operatorname{Hopf}}(X)=\mathcal{A}(\gamma)$ and ${\rm CZ}(\gamma)=3$. It follows from equation (3.3) that the ECH index $I(\gamma)=2$.
Now choose a generic almost complex structure $J$ as needed to define the ECH differential $\partial_{J}$, and choose a point $y\in Y$ as needed to define the chain map $U_{J,y}$. We claim that $\partial_{J}\gamma=0$ and $U_{J,y}\gamma=\emptyset$. Assuming this claim, it follows from the nontriviality of the class []delimited-[][$\emptyset$][ ∅ ] that $\gamma$ is a cycle in the ECH chain complex representing the generator of $\operatorname{ECH}_{2}(Y,\xi_{0})$, so it follows from Definition 3.8 that $c_{1}^{\operatorname{ECH}}(X)\leq\mathcal{A}(\gamma)$, completing the proof of Step 1.
To prepare to prove the claim, recall from Remark 1.9 that since $X$ is dynamically convex and $\gamma$ is a Hopf orbit, it follows that $\gamma$ is the boundary of a disk-like global surface of section $D\subset Y$. In particular, every Reeb orbit other than $\gamma$ has positive linking number with $\gamma$. Moreover, the construction of the global surface of section in [hryn_jsg] obtains $\operatorname{int}(D)$ as the projection to $Y$ of a $J$-holomorphic plane in ×$\mathbb{R}\times Y$ with a positive puncture at $\gamma$.
To prove that $\partial_{J}\gamma=0$, we will show that if $\beta$ is an ECH generator with $I(\beta)=1$, then $\mathcal{M}^{J}(\gamma,\beta)=\emptyset$. Suppose to the contrary that there exists $C\in\mathcal{M}^{J}(\gamma,\beta)$. We can write $\beta=\beta_{1}^{m_{1}}\cdots\beta_{k}^{m_{k}}$. Since $I(\beta)\neq 0$, we must have $k>0$. By the inequality (3.5), all of the orbits $\beta_{i}$ are distinct from $\gamma$.
It follows from the asymptotics in Theorem 2.4 that if $a_{0}$ is sufficiently large, then $C$ is transverse to {±$\{\pm a_{0}\}\times Y$, and
$C\cap([a_{0},\infty)\times\gamma)=C\cap((-\infty,a_{0}]\times\gamma)=\emptyset.$
Let $\eta_{\pm}$ denote the projections of $C\cap(\{\pm a_{0}\}\times Y)$ to $Y$. We have
(3.6) $\mathrm{link}(\eta_{+},\gamma)-\mathrm{link}(\eta_{-},\gamma)=\#(C\cap([-a_{0},a_{0}]\times\gamma))\geq 0,$
where ##\## denotes the algebraic count of intersections, which is nonnegative by intersection positivity of holomorphic curves.
Let $\tau$ be a global trivialization of the contact structure $\xi$, and let $\nu$ denote the asymptotic eigenvalue of $u$ at $\gamma$ as in Theorem 2.4. Since $\nu<0$ and $\operatorname{CZ}_{\tau}(\gamma)=3$, it follows from the discussion in §2.2 that
$\operatorname{wind}_{\tau}(\nu)\leq 1.$
On the other hand, it follows from the definition of self-linking number that
$\operatorname{link}(\eta_{+},\gamma)=\operatorname{wind}_{\tau}(\nu)+\operatorname{sl}(\gamma).$
Since $\operatorname{sl}(\gamma)=-1$, we conclude that
(3.7) $\operatorname{link}(\eta_{+},\gamma)\leq 0.$
On the other hand, we have
(3.8) $\operatorname{link}(\eta_{-},\gamma)=\sum_{i=1}^{k}m_{i}\operatorname{link}(\beta_{i},\gamma)\geq k\geq 1.$
Combining the inequalities (3.6), (3.7), and (3.8) gives $-1\geq 0$, a contradiction.
To prove that $U_{J,y}\gamma=\emptyset$, we can choose $y\in\operatorname{int}(D)$. Let $C\in\mathcal{M}^{J}(\gamma,\emptyset)$ be a $J$-holomorphic plane that projects to $\operatorname{int}(D)$. By translating $C$ in the $\mathbb{R}$ direction, we can assume that $(0,y)\in C$. We claim now that if $\beta$ is an ECH generator, if $C^{\prime}\in\mathcal{M}^{J}(\gamma,\beta)$, and if $C\neq C^{\prime}$, then $C\cap C^{\prime}=\emptyset$, and in particular $(0,y)\notin C^{\prime}$. This will prove that $C$ is the only curve counted by $U_{J,y}\gamma$ so that $U_{J,y}\gamma=\emptyset$.
To prove that $C\cap C^{\prime}=\emptyset$, observe that by the definition of the relative intersection number in [ind_rev, §2.7], we have
(3.9) $Q_{\tau}(\gamma)=\#(C\cap C^{\prime})-\ell_{\tau}(C,C^{\prime}),$
where $\ell_{\tau}(C,C^{\prime})$ is the asymptotic linking number of $C$ and $C^{\prime}$, which is the linking of the positive ends of $C$ and $C^{\prime}$ in a neighborhood of $\gamma$ with respect to the trivialization $\tau$; see [ind_rev, Def. 2.9] for the precise definition. Since $\tau$ is a global trivialization and $\operatorname{sl}(\gamma)=-1$, it follows from equation (3.2) that
(3.10) $Q_{\tau}(\gamma)=-1.$
On the other hand, it follows from the linking bound in [ind_rev, Lem. 4.17] that
(3.11) $\ell_{\tau}(C,C^{\prime})\leq 1.$
Combining (3.9), (3.10), and (3.11), we obtain #$\#(C\cap C^{\prime})\leq 0$. By intersection positivity, it follows that #$\#(C\cap C^{\prime})=0$ and $C\cap C^{\prime}=\emptyset$ as desired.
Step 2
We now show that any cycle in the chain complex $(ECC_{*}(Y,\lambda),\partial_{J})$ that is homologous to $\emptyset$ must include $\emptyset$ as a summand. For this purpose it suffices to show that if $\alpha$ is an ECH generator with $I(\alpha)=1$, then $\mathcal{M}^{J}(\alpha,\emptyset)=\emptyset$. One can prove this using the Fredholm index, but to prepare for the following step we will give a different proof using the “$J_{0}$ index” defined in [ind_rev, §6.1].
In the present context, if $\alpha=\prod_{i=1}^{k}\alpha_{i}^{m_{i}}$ is an ECH generator, then $J_{0}(\alpha)$ is an integer defined by
(3.12) $J_{0}(\alpha)=I(\alpha)-\sum_{i=1}^{k}{\rm CZ}(\alpha_{i}^{m_{i}}).$
The key property of $J_{0}$ that we need is that if $\beta$ is another ECH generator with $I(\alpha)-I(\beta)\in\{1,2\}$, and if there exists $\mathcal{C}\in\mathcal{M}^{J}(\alpha,\beta)\neq\emptyset$ (i.e. if $\mathcal{C}$ is counted by the ECH differential or the $U$ map from $\alpha$ to $\beta$), then
(3.13) $J_{0}(\alpha)-J_{0}(\beta)\geq-1,$
with equality only if $\mathcal{C}$ consists of a holomorphic plane, possibly together with some trivial cylinders. This is a special case of [26, Prop. 5.8].
Suppose now that $\alpha=\prod_{i=1}^{k}\alpha_{i}^{m_{i}}$ is an ECH generator with $I(\alpha)=1$. Since $I(\alpha)\neq 0$, we must have $k\geq 1$. It then follows from equation (3.12) and dynamical convexity that
$J_{0}(\alpha)\leq-2.$
On the other hand $J_{0}(\emptyset)=0$. Then if $\mathcal{M}^{J}(\alpha,\emptyset)\neq\emptyset$, the inequality (3.13) with $\beta=\emptyset$ gives $-2\geq-1$, a contradiction.
Step 3
We now prove that $c_{\rm Hopf}(X)\leq c_{1}^{{\rm ECH}}(X)$.
By Definition 3.8, there is a cycle $\sum_{j}x_{j}$ in the ECH chain complex $(\operatorname{ECC}_{*}(Y,\lambda),\partial_{J})$, representing the generator of $\operatorname{ECH}_{2}(Y,\xi_{0})$, such that each summand $x_{j}$ is an ECH generator with ECH index $I(x_{j})=2$ and action $\mathcal{A}(x_{j})\leq c_{1}^{\operatorname{ECH}}(X)$. Since $U:\operatorname{ECH}_{2}(Y,\xi_{0})\to\operatorname{ECH}_{0}(Y,\xi_{0})$ is an isomorphism, and since $\emptyset$ is a cycle representing the generator of $\operatorname{ECH}_{0}(Y,\xi_{0})$, it follows that $U_{J,y}\sum_{j}x_{j}$ is homologous to $\emptyset$. By Step 2, there is some $j$ such that $\emptyset$ is a summand of $U_{J,y}x_{j}$, and in particular there exists $\mathcal{C}\in\mathcal{M}^{J}(x_{j},\emptyset)$.
By the inequality (3.13), we have $J_{0}(x_{j})\geq-1$. By equation (3.12) and dynamical convexity, $x_{j}$ consists of a single simple Reeb orbit $\gamma$ with multiplicity $1$, and $\operatorname{CZ}(\gamma)=3$. Since $I(\gamma)=2$, it follows from equation (3.3) that $\operatorname{sl}(\gamma)=-1$. Since equality holds in (3.13), $\mathcal{C}$ is a holomorphic plane positively asymptotic to $\gamma$. By [30, Prop. 3.4], the projection of $\mathcal{C}$ to $Y$ is an embedding [^9] The statement of [30, Prop. 3.4] requires that $\gamma$ is elliptic, but the proof works just as well if $\gamma$ is negative hyperbolic, so it applies given that $\operatorname{CZ}(\gamma)=3$. The basic ideas in the proof go back to [props2], and a more general statement is proved in [sief_int, Theorem 2.6]. One can further show that the projection of $\mathcal{C}$ to $Y$ is the interior of a global surface of section for the Reeb flow, but we do not need this.. Thus $\gamma$ is unknotted in $Y$, so $\gamma$ is a Hopf orbit. Then $c_{\operatorname{Hopf}}(X)\leq\mathcal{A}(\gamma)$, and we know from the previous paragraph that $\mathcal{A}(\gamma)\leq c_{1}^{\operatorname{ECH}}(X)$.